OpenStudy (anonymous):

If f(x) = log2 (x + 4), what is f−1(3)?

OpenStudy (anonymous):

Given the function f(x) = 0.5(3)x, what is the value of f−1(7)? 0.565 1.140 1.771 2.402

OpenStudy (ranga):

If "f-1" the inverse function in the first problem, then to find f^-1(3) you need to set the left hand side to 3 and solve for x. f(x) = log2 (x + 4) 3 = log2 (x + 4) (x+4) = 2^3 = 8 x = 8 - 4 = 4 So f^-1(3) = 4.

OpenStudy (anonymous):

the x next to the (3) is an exponent

OpenStudy (anonymous):

Given the function f(x) = 0.5(3)x, what is the value of f−1(7)? 0.565 1.140 1.771 2.402

OpenStudy (ranga):

f(x) = 0.5(3)^x, what is the value of f−1(7)? 7 = 0.5 * (3)^x 7/.5 = (3)^x 14 = 3^x log(14) = x * log(3) x = log(14) / log(3) = 2.402

OpenStudy (anonymous):

I figured the other one out can you help with this one

OpenStudy (anonymous):

can you help me with 2 more

OpenStudy (anonymous):

Given the parent functions f(x) = log10 x and g(x) = 5x − 2, what is f(x) • g(x)? f(x) • g(x) = log10 x5x − 2 f(x) • g(x) = log10 (5x − 2)x f(x) • g(x) = 5x log10 x + 2 log10 x f(x) • g(x) = 2 log10 x − 5x log10 x